IR-Master

An Infrared Control Device For Amiga Computers Version 3.11

January 1995

by Juergen Frank and Michael Watzl

Copyrigth \begin{matrix}c \\ ◯ \end{matrix} 1993-1995 by Jürgen Frank and Michael Watzl

1 What is IR-Master

The IR-Master project consists of two software components and a control hardware, which is connected to the joystick port of your Amiga (masterhardware). With this you can control almost every device which is supplied with a ir remote.

At first you start edition your virtual remote control with the IR-Editor. This one is saved to disk and activated with the IR-Runner. The number of IR-Runners used at one time is not limited.

Your edited remote control opens as a window on the workbench. With this window you can interactively control your device[s], you can use the IR-Runner’s ARexx-port for programming your devices, See Section 4.3 [ARexx Port], page 12. Finally you can send infra red codes from within other applications or edit scripts which will automatically control e.g. your video recorder.

With the version 3.0 of the IR-Master two new hardware components are introduced. Both are optional. The first is designed for the parallel port and consists of up to 8 ir senders which can be accessed separately. The other hardware is designed for the serial port and uses DMA to send the ir signals.

Starting with version 3.1 the serial hardware can also be connected to the left and right audio channels of the amiga. The output channel may be selected for each gadget.

However, in any case you need the hardware connected to the joystickport for learning the ir commands.

Since version 2.5 of the IR-Master two new features were introduced: You can directly send commands to LControl and Rapid. This allows you to control video cameras supported with an LANC/Control-Linterface and video recorders supported with an GSE-Rapid-interface. For more information, please consult Section C.4 [LControl & Rapid], page 22.

2 Installation

Requirements:

req.library

may not be includet in this package, but can be easily found - it’s freely distribatable and written by Colin Fox (Pyramid Designs) and Bruce Dawson (Cygnus Software), See Section C.1 [req.library], page 20.

Kick/WB 1.x

this version in no more compatible to the harware. It is therefore no more supported.

Kick/WB 2.x/3.x

for the fine and new (also working :-)version.

Installation:

1. Copy the ‘req.library’ to your ‘LIBS:’ drawer of your boot disk.

A patch (like RTPatch from Nico Francois) is recommended to patch the req.library calls to the much better looking reqtools.library (also from Nico Francois ->thanks). See Appendix C [Other Software], page 20.

2. Copy tht ‘InfraRed.library’ to the ‘LIBS:’ drawer

3. Create a directory for the IR-Editor and the IR-Runner and copy them to this directory. However, it is not needed for the two programs to be in a certain drawer. No assigns are needed. Note: Your IR-Runner should be placed in a directory that is in the workbench search path (which is declared in your ‘startup-sequence’ right before ‘loadwb’), e.g. ‘C:’ or ‘sys:utilities’. So the projects will surely find the default tool.

4. If you use WB 2.1 or higher:

Copy the languages-catalog you want to use from ‘catalogs/<language>’ to ‘LOCALE:catalogs/<language>’. At the moment there is only german, danish and english (built-in).

3 IR-Editor

With the IR-Editor you can create a remote control on your Amiga screen.

3.1 The Menus

3.1.1 Project

Load... Pops up a requester to load a saved project. Save Saves the current set of gadgets.

Save as...

Pops up a requester to save the current project.

——————

Info Gives some information about the used memory etc. About Credits...

——————

Save Prefs

Saves the current settings, like position of the toolbox, size of project window...

Iconify Closes screen, frees memory and puts an app icon onto the workbench

——————-

Quit Quits the editor.

3.1.2 Edit

Cut Cuts the selected Gadgets to a internal buffer. That means they will be deleted in the

projectwindow and they can be pasted back anywehere.

Copy Copies the selected Gadgets to a internal buffer. With "paste" you can insert this gadgets whereever you want. Note: If the buffer is not empty and you cut/copy gadgets than the gadgets already in the buffer will be overwritten!

Paste Paste back gadgets in the internal buffer.

Info Gives an information of the gadgets which are stored in the buffer.

Flush Clears the internal buffer.

3.1.3 Arrange

Note: The reference gadget is always the first selected. That means, if I first select a big gad and than two small gads and I select clone size then all gadgets will be big...

Align Gadgets

aligns the selected gadgets

Left by the left edge or

Top by the top edge

Spread Gadgets

Horizontally pops up a requester which asks how many pixels

Vertically should be placed between each of the selected gadgets in the selected di-

rection...

Clone Size

Clones the sizes of the selected gadgets.

——————-

Coordinates

Opens/Closes the coordinates window. Only if one (!) gadget is selected the coordinates will be shown an can be directly edited. ——————-

Grid Activates/inactivates the grid for better and easier placing and arranging of the gadgets

Adjust Grid

Here you can select the grid size

3.1.4 Environment

Palette Pops up the standard palette requester.

Font Brings a font requester to the front. The selected font will only be used for the editor’s screen. To change the font of the gadgets you have to use the font gadget in the toolbox.

Screenmode

Pops up a screenmode requester.

On Public Screen

Tries to close the IRMaster screen and reopen on the default public screen (mostly the wb).

Make Screen Public

Tries to close the current IRMaster GUI and reopen it as a public screen. The screen’s name is "IRMaster".

Show IR-Code

Shows the IR-Code of the current gadget (if learned). It is designed like an oszillator, so you can easyly decide if the last learn action was successful! Length specifies the amount of memory which is used to send this command. The percentage shows how much of the memory reserved for this command is used.

Show Configuration

Shows the IR-configuration of the current gadet. For more information about the ir-configuration, See Section 3.4 [Learning IR-Commands], page 9.

3.1.5 Hardware

With the version 3.0 of the IR-Master the joystickport-hardware was changed. To remain compatible to the old hardware this menu has been implemented.

Choose "Old" if you own a hardware designed for lower versions than 3.0.

3.2 The Toolbox

3.2.1 NewGadget

shortcut F1

This function brings you to the "new gadget mode". You can see a border attached to your pointer. With a mouseclick you can place the new gadget. A new gadget is always named ‘New’ and its size is 50 x 20 pixels.

3.2.2 Move Gadget(s)

shortcut F2 If you have some gadgets selected and you press the move button or the move shortcut the

selected gadgets will be represented by borders which stick to your mouse pointer. With a click you can put the gadgets to their new position.

3.2.3 Attache Grafic To Gadget

shortcut F3 For each of the selected gadget a file requester pops to the front, asking you for an icon/brush.

This image will be used as gadget imaginary.

3.2.4 Attache Text to Gadget

shortcut F4 Pops a text requester to the front (for each gadget). You can type any text. The length is not

checked! It doesn’t have any serious consequences if the text is to wide for the gadget - it simply looks awful.

3.2.5 Learn IR data

shortcut F5 With this gadget you can switch to the learn mode. The pointer will remind you being in this

mode. If you click a gadget, the Amiga is frozen and waits for ir input to the IR-Sensor, see also Section 3.4 [Learning IR-Commands], page 9.

3.2.6 Kill Gadget(s)

shortcut F6

Removes all selected gadgets without warning! They are unrecoverably lost.

3.2.7 Size Gadge

shortcut F7 Switches to the size mode. A border is drawn - the lower right corner sticks to the pointer.

After a mouseclick the gadget will be resized and the text/grafic will be centered in the gadget.

3.2.8 Load Backgound Grafic

shortcut F8 Pops up a requester. Here you can select any ‘.ilbm’-picture to be used as background grafik

for your remote control. This grafik’s name and path will be set in the tooltypes of the project when it is saved.

Therefore the grafic must exist when the project is run. The palette of the picture won’t be used! I’m sure you like your wb-colors more!

To clear the background again, click this gadget (or press F8) and select cancel in the filerequester.

3.2.9 Change Font

shortcut F9 Pops up a font requester. The selected font will be used for all the selected gadgets. Make sure

that this font is also available to the IR Runner.

3.3 Test IR Data

shortcut F10

Switches to the test mode. The pointer reminds you being in the test mode. Note: To test ir commands, they must be learned before. To test LControl or Rapid commands, LControl/Rapid must be running. See Section C.4 [LControl & Rapid], page 22.

3.4 Learning IR-Commands

Learn ir-commands:

• toggle to learn mode by pressing F5 or the equivalent gadget

• click on the gadget whose function you want to learn ↦ pointer is frozen...

• point your remote control to the ir-sensor (distance ca. 1-15 cm)

• hold the button down ’till the pointer can be moved again

Test ir-commands

• switch to test mode by pressing F10 or the equivalent gadget

• hold transmitter in direction to the ir-receiver of the implement with a distance of 1-30 cm.

• click on the gadget you want to test ↦ command will be sent

Important notes

1. It is possible that the implement (like CD, Videorecorder, TV set,…) won’t react on the irmaster the first time. In this case you have to change the parameters in the config-window, See Section 3.1.5 [The Menus], page 5. To get to the configrequester you must:

• click on the gadget the config should be changed

• open the config window All the changes (but not wobbelcode) can be testet without learning it again. Config menu:

- Configuration

37 kHz Carrier frequency of the ir signal is 37 kHz, matches for ca. 70% of all imple-

ments

45 kHz Carrier frequency is 45 kHz, matches for ca. 10%

66 kHz Carrier frequency is 66 kHz, matches for some implements

Unpulsed no carrier frequency, ca. 8%

LControl Directly sends commands to LControl to control video cameras via LANC, See Section C.4 [LControl & Rapid], page 22.

Rapid Directly sends commands to Rapid to control video recorders via GSE-Rapid-Code, See Section C.4 [LControl & Rapid], page 22.

Command Executes a DOS-Command. The text is interpreted as a line in the shell. Relais-Card

The IR-Master is also able to control a relai-card (special support for the relai-card which can be accessed by Konrad Electronic, Hirschau, Germany, art-nr: 192252-55 + 174025-55). This is a 4 channel ir receiver. With this receiver you can control any devices (like lamps, motors, etc.). Die ir-codes are already learded in by the IR-Master, so you do not have to relearn them. There are 2 number fields: The "RL:"- and "CH:"-field. The first is the bitcoded address of the relai(s) you want to activate and the second field is for the bitcoded address for the channel(s) you want to access. The "address" of the relai is 4 bit wide and the "address" of the channel is 6 bit wide. Each relai is represented by one bit (also the channels). Example

• If you want to activate the relais 2 and 3 on channel 0 you have to type the value of binary %1100 (which is 12) into the relai field and 0 into the channel field.

• If you want to activate relai 1 on the channels 2,3 and 4, you have to type 1 into the relai field and the value of %11100 (which is 28) into the channel field.

DMA-Audio-Send If selected the ir singal is send to the (optional) hardware connected to the serial port or the audio channels of the amiga. DMA is used and therefore the pointer will not be frozen for a short moment. Please read the description of the hardware for more info. The polarisation (positive/negative) must be choosen correctly, too. This is best done by trying it out. If you have a two-channel audio hardware you may also select the output channel (left or right).

- Send command slider This indicates how often the ir command should be sent. Normally 2-5 times is enough (the lower it works the better it is).

- Wobbelcode This is a special form of ir code. An implement which uses Wobbelcode (e.g. Grundig tools) sends code A the first time and the next time it sends (for the same function) code B. In this case you have to learn twice. To learn this type of code you must click on Wobbelcode in the config window and then click on learn and follow the messages in the status bar. You can recognize Wobbelcode-implements when the implement reacts the first time at once but the second time it does nothing or it reacts very late.

- IR-Channel This configuration is only for those who own a hardware for the parallel port. Other user should always have onyl the M activated. "M" stands for master hardware which is the one connected to the joystickport. The other gadgetst (1-8) represent the senders on the parallel hardware. So you can send a signal to channel 1,3 and the joystickport hardware by simply selecting them.

2. The learning of ir commands depends on the distance between remote control and ir sensor. After a faulty try you should attempt it again with different distances between 1 and 20 cm.

3. If the batteries of the remote control are weak you should use new batteries.

4. In general all impements can be controlled exept Telefunken. For this type of implements we offer a special hardware - simply contact us and describe your implement.

5. Finally: The parameters in the config window are the most common but if it does not work - use trial and error. You can’t damage your hardware by changing the parameters... believe me ;-)

If you own a Akai, Blaupunkt, Fisher, Technics, Philips, Grundig, Panasonic, Sony... you can learn the command set by nearly 100%. (tested very (!) often...)

You can also use up to 5 (with y-adapters) ir transmitter in your joystickport.

3.5 Drag & Drop

‘Drag & Drop’ is a Feature of AmigaDOS 2.04 and above which enables you to drag an icon and drop it onto a special window causing a special action. This is - of course - only possible when the applikation has windows opened on the workbench. To open your ir-editor on the workbench, See Section 3.1.5 [The Menus], page 5.

This feature is implemented in the following way:

You can drag an iff-ilbm-picture onto the project/toolbox-window. This picture will be loaded as background grafik. You can also drop a project’s icon onto the window. Then, this project will be loaded at once.

The file type recognition works via the file name. If the name’s suffix is ‘.iff’, ‘.ilbm’ or ‘.brush’ the editor assumes a picture file. If the suffix is ‘.irm’ the editor tries to load a project.

4 IR-Runner

The IR-Runner is used to start finished projects. Multiple IR-Runners may be in use.

4.1 Starting The IR-Runner

There are different possibilities to start a project:

1. Shell/Cli

         1> IRRunner <project.irm> LEFT=20 TOP=20 PUBSCREEN=TERM

2. Workbench • Click on IRRunner (once) + shift-doubleclick on the project’s icon

• Make sure that the IR Runner is in the workench searchpath (the path set up before

        ‘loadwb’ is called in the ‘startup-sequence’, e.g. ‘C:’, ‘sys:utiltiies’)

Doubleclick the project’s icon. (default tool must be correctly set)

• If you doubleclick the ir-runner a filerequester will appear...

4.2 IR-Runner Menus

Each remote control has its own gui and its own menu:

About Gives you VERY important information about you, god and the universe. Hide Closes the gui. The arexx port will still be available.

Fix Position

Saves the current IR-Runner window position to the project’s icon (as tooltypes).

Quit Quits this remote control.

4.3 ARexx Port

The IR-Runner has it’s own rexx port. The port name depends on the project’s name:

Project name: CD_Player.irm ↦ port name: CD_Player_rexx

Now the commands:

IR_Send The IR Send command needs one parameter:

• Either the gadget number (the number shown in the editor when the gadget is selected)

• or: the gadget’s text (if it is no grafic!) Therefore you must specify a $ followed by the gadget text.

From the shell this could look like this:

     1> rx "address 'CD_Player_rexx' IR_Send 3"

or in a rexx script:

     /* cd player -> play */
     address 'CD_Player_rexx'
     IR_Send '$Play'     /* assuming that the gadget's text is 'Play' */
     exit

IR_Move With the IR Move command you can move the window (if opened). This command needs two arguments, the x and y coordinate of the new window position. Example:

     IR_Move 100 120

IR_Quit Quits the application.

IR_Disappear

Closes the gui without closing the rexx port. It acts like selecting hide from the menu/clicking hide in commodities exchange. No arguments are needed.

IR_Appear

Opens the gui (if cosed) and (if cosed) reopens the rexx port. No args.

IR_Disable

Acts like IR Disappear but the rexx port is closed, too. More memory is freed again. No args.

IR_Enable

Reopens (if closed) the rexx port and reopens (if closed) the gui. No args.

4.4 Tooltypes

The IR-Runner recognizes the following tooltypes, when given in the project icon:

CX_POPKEY="string"

Here you can specify a hotkey - with this shortcut, you can activate this IR-Runner project Example: CX_POPKEY=lalt f10 If you hold left-alt and press f10, this IR-Runner project wil be activatet.

CX_POPUP=YES|NO

If you specify NO, no window will be opened when starting this project. You can control this project only via ARexx and Commodities-Exchange. However you can open the project’s window by ARexx/Cx.

LEFT=xxx Here you can specify the default left coordinate of the window. This tooltype is automatically generated when selecting fix position in the menu.

TOP=yyy The default top coordinate of the window.

BACKGROUND=name

Here the background grafik is specified.

PUBSCREEN=name

name is the title of a public screen on which the IR-Runner should open its gui.

HARDWAREVERSION=

If you own a hardware which is designed for an IR-Master version lower than 3.0 you should set this value to 0. Otherwise it should be set to 1. However, this tooltype is set by the IR-Editor automatically.

5 Hardware

With the version 3.0 of the IR-Master a new hardware is introduced. The new hardware is more powerful than the old one. To remain compatible to the old hardware you can select the hardwaretype in the IR-Editor and the IR-Runner.

Use the menus in the IR-Editor and the tooltypes in the IR-Runner to select the old hardware. See Section 3.1.5 [The Menus], page 5. See Section 4.4 [Tooltypes], page 13.

5.1 Update old Hardware

To update your old hardware (turn IR-Master-1.0/2.61-hardware into a IR-Master3.0-hardware) you must exchange T1 into a BC 560 or BC 559.

Alternatively you can also rebuilt everything corresponing to the new hardware schematics.

After you updated the hardware it will have a range of 4-6 m.

Attention: Do not use an old hardware and a new hardware at one time! When you updated the hardware you must use IR-Master 3.0 to avoid overloading the new hardware.

See chapter IR-Editor/Section 3.1.5 [The Menus], page 5, look for "hardware" and chapter IR-Runner/Section 4.4 [Tooltypes], page 13, look for "HARDWAREVERSION=".

5.2 Tuning

The normal range of the IR-Master 1.0 to 2.61 hardware is ca. 1-2 m. The new one reaches already 4-6 m.

Both types of the hardware can be tuned. But before you tune your old hardware you should think about updating it. See Section 5.1 [Update old Hardware], page 15.

5.2.1 Tuning the old Hardware

There are several possible modifications:

• exchange R2 with 150 Ohm; range: about 3 m

• exchange R2 with 100-120 Ohm; range about 4 m X)

• 2 x D1 in row, R2:100 Ohm; range about 5 m

• 2 x D1 in row, R2:82 Ohm; range about 6 m X)

• exchange D1 with LD 273 or SFH 485 (e.g. Konrad Elec.) and R2: 75 - 100 Ohm; range: up to 10 m Important: The marked (x) tunings need special care for the hardware! The hardware must only be plugged

in when a IR-Runner/IR-Editor is running. Otherwise the hardware will be destroyed after a while. Remark: Before you tune your old hardware you should better think about making a new hardware out

of your old hardware. The new hardware has a range of 4-5m.

5.2.2 Tuning the new Hardware

The new hardware (IR-Master 3.0) already has a range of 4-6 m. If this is not enough for your needs you can modify your hardware in the following ways:

• exchange R2 to 47 Ohm; range: about 8 m; low scattering/spreading

• 2 x D1 in row and R2 = 39 Ohm; range: over 10 m; good scattering/spreading

• 3 x D1 in row and R2 = 18 Ohm; range: up to 18 m; very good scattering/spreading

5.3 Joystickport-Hardware

With the hardware connected to the joystickport ir-commands can be learned and sended. This hardware is obligatory, because the other hardwares are send-only.

If you own a IR-Slave, you can use this hardware for learning, too. The results are even better.

5.4 Parallelport-Hardware

The hardware for the parallelport is "send-only". It has up to 8 senders. For each gadget/ir command you can select which senders should be activated when you click it.

Now it is possible to control two or more devices of the same type. You can for example control device 1 via the sender 1 of the parallelport-hardware and the device 2 via the joystickport-hardware.

5.5 DMA/Audio-Hardware

The DMA/audio-hardware uses DMA when sending. The Amigas multitasking need not to be stopped for a moment (like it must be done for the other hardwares).

It uses the left audio channel for sending. The Amigas serial port supplies the neccessary voltage and the audio signals, too. Therefore the serial port is used for this hardware.

Beginning with version 3.1 a 2 channel hardware is supported. You may now select The left or right channel for sending.

For experts it should be no problem to built a hardware which will be connected to the left audio channel with an external power supply. The serial port remains free, then.

Appendix A Support & Bugreports

A.1 Support

You can send a disk + enough stamps to the following adresses to get the newest version of IR-Master:

Michael Watzl               or      Juergen Frank
Konradstr. 11                       Wittelsbacherweg 7
86609 Donauwoerth                   86609 Donauwoerth
Tel: +49 906 5834                   Tel: +49 906 1057
Germany                             Germany

The new version will be immediately sent back.

A.2 Bugreports

Make sure, that the bugreport descibes the bug AND your configuration as well as possible. If possible describe the very exact way how to produce the mentioned bug. Bugs are so much

easier to find if you can reproduce them...

Appendix B Registration

REGISTRATION FORM (3.11)
NAME:       ___________________________________________
STREET:     ___________________________________________
ZIP/CITY:   _________________________________________
COUNTRY:    ___________________________________________
TELEPHONE:  _____________
I want to become registered user of the IRMaster. Therefore I
o enclose a cheque
(foreign countries: drawn on a german bank
europe           : euro cheque)
o enclose cash
Please send me:
o newest version of this package + hardware schematics      $15
o newset version of this package + constr. kit + schematics $25
o newest version of this package + compl. hardware (+ docs) $40
----- optional hardware ------------
sender modules for the parallel port (incl. one sender)
constr. kit:
o sender modules for the parallel port (incl. 1 sender)  $6
o __x additional sender (up to 7 additional ones)       $1
ready to use:
o sender module for the parallel port (incl. 1 sender)   $8
o __x additional sender (up to 8 additional ones)       $2
-------------------------------------
The complete order will make ______$.
_______________________________________________
   date, locality                            sign

Appendix C Other Software

C.1 req.library

Taken from "req.doc" (description on req.library):

"The req.library is a run time re-entrant library that is designed to make it easier for programmers to use powerful, easy to use requesters for communicating with users. The requester library includes such functions as a color requester, file requester, message display requester and many functions to make the creation of gadgets for your own custom requesters easier.

Req.library was written by Colin Fox (of Pyramyd Designs) and Bruce Dawson (of CygnusSoft Software). Req.library is a freely distributable library that may be used in commercial products without paying any royalties. We encourage you to use the requester library in all of your programs, to make them easier to write, and to use.

Req.library is not public domain. The requester library and all documentation and example programs are all copyright 1989."

C.2 RTPatch

Taken from RTPatch.doc:

"The programs and files in this distribution are freely distributable, but are also copyright (c) Nico François. They may be freely distributed as long as no more than a nominal fee is charged to cover time and copying costs. No commercial usage is permitted without written permission from the author. Everything in this distribution must be kept together, in original unmodified form. The above is generally known as freeware.

If you have suggestions or remarks about this program, or if you find any bugs, please let me know.

        Contacting the author:
          Fido:     2:292/603.10 (Nico Francois)
          Internet: nico@augfl.be
          Mail:     Nico François
            Corbielaan 13
            B-3060 Bertem
            BELGIUM

If you can please use e-mail. That way you’ll stand a much better chance of getting a reply quickly."

C.3 reqtools.library

Taken from reqtools-distribution’s readme file:

"reqtools.library is a standard Amiga shared, runtime library. The purpose of ReqTools is to make it a lot quicker and easier to build standard requesters into your programs. ReqTools is designed with Commodore’s style guidelines in mind, all requesters have the look-and-feel of AmigaDOS Release 2.

If you use ReqTools in your programs you do not have to worry about requesters too much. You’ll have a lot more time to concentrate on the really important stuff your program does.

The library has been inspired by req.library (by Colin Fox & Bruce Dawson) and Commodore’s own asl.library. You will find it shares a lot of the features with these two libraries, while it has some additional ones as well. While asl.library was great in look-and-feel I felt it lacked quite a bit in the power department. This power is mostly available in req.library, but this doesn’t fit in too well with AmigaDOS Release 2. So the idea for a new requester library was born.

I tried to make all requesters in ReqTools very powerful, easy to use and with a very nice new-look 3D interface. Efficiency was also considered very important. The 2.0 version of the library is about 38K large. I’m sure you will agree this is not too much considering what ReqTools offers.

ReqTools features the following requesters: a query requester, a string requester, a number requester, a file requester, a font requester, a palette requester and new in release 2.0 of ReqTools a volume requester and a screen mode requester."

C.4 LControl & Rapid

LControl and Rapid are other programs by Juergen Frank and Michael Watzl.

LControl Lcontrol is a program which allows you to control a video camera (supportet with a LANC/Control-L interface). The control hardware is rather simple to construct. The hardware is connected to the parallel port of the Amiga.

LControl has an ARexx port and is a commodity.

Rapid If you own a video recorder which is supplied with an GSE-Rapid interface than you can use your Amiga and Rapid to control it! At the momemt rapid is not yet finished (it is still in the ß-phase). If you want to know more about the Rapid-project contact the authors.

Rapid uses the serial port.

C.5 IR-Master History

V1.0
    First Released Version  (version for OS 1.3)
V2.0
    Big rewirte for 2.x/3.x
    Many features added:
- arrangement funtions
- environment menu
- font support for gadgets
V2.01
    Bugs fixed:
- OS3.0 menu new look added
- several little bugs removed
    Using asl requesters in stead of req requesters
    req.library is only used for GetString();
V2.02
    Bug fixed:
- Coords used system default font in stead of topaz 8
    Features added:

- Screen can be made public
V2.05
    Some minor bugfixes...
    Features added:
- locale is supported + german cataloges added
- can open on a public screen
V2.06
    IR_Runner features added:
- Now supports fixing window position
    IR_Runner bugs removed:
- No more gurus when sending IR commands
V2.07
    IR_Runner bug removed:ed:ion for OS 1.
- Big stack bug removed!
    IR_Editor
- minor bug in "project-changed" removed
V2.07+
    Docs: Harware tuning infos added
V2.08
    IR-Editor
- IR-Code window added
- Brushs can be used for gadget imaginary
    IR-Runner
- Arexxport improved
V2.09
    IR-Editor
- Background gfx now supported!!!
- IR-Config window redone
V2.1b
    IR-Editor
- new icons in toolbox :-)
- bugs in asm-code removed
    IR-Runner
- better handling of multiple runners
- tooltype handling improved
V2.11
    IR-Editorinfos added inf
- Drag & Drop added
- better tooltype handling
    IR-Runner
- Bug in Filerequester removed

- Arexx-port improved
  See Doc: IR_Send for importand changes!
- better tooltype handling
V2.2
    IR-Editor
- Cut/Copy/Paste added
    Docs:
- DVI/TeX-Documentation added (only german so far)
V2.21
    IR-Editor
- Bug in wobbelcode learn removed
    IR-Runner
- Bug in wobbelcode send removed
V2.22
    IR-Editor
- PubScreen Name is now "IRMaster"
V2.50
    IR-Editor/IR-Runner
- New Gadgettypes:
    > Direkt LCOntrol and Rapid support
    > DOS/ARexx-Kommands on Gadgets
    IR-Editor
- Bugs in "Paste" removed
    IR-Runner
- Enforcerhits detected and traced down...
V2.60
    IR-Editor/IR-Runner
- New Gadgettype: Relais-Card
  This type has a already lernd in systemcode to control
  the Relais-Card by Conrad-Electronics [for more, read the
  Docs]
- Bug removed: If you use different Fonts in the IR-Editor
  & Runner, the display now won't get confused.
V2.61
    Minor Bugfixes in the Config-Window
V3.00
    IR-Editor/IR-Runner
- Supports now special hardware for the
  parallel and serial port
- Raster in IR-Editor
- Faster signal repetition when holding down a
  button in the runner
- several bugfixes

- changed hardware !!!! improves overall performance
  new ir-range: 5-6m!
V3.10
    IR-Editor
- better handling of gadgets (stay selected...)
    IR-Editor/IR-Runner
- supports now colored text gadgets
- audio-hardware supports left and right channel
  (like the parallel hardware)
V3.11
    IR-Editor and IR-Runner
- bugfix in DMA-send routine removed
- minor bugfixes in InfraRed.library
  (current version: 35.5)

Index 26

Index

A

About. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 12 Addresses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ARexx Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Arrange . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Attache grafic to gadget. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Attache text to gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Authors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

B

Background grafic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Background picture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 BACKGROUND= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 BBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Bruce Dawson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 brush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

C

Change Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Colin Fox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Control-L. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 CX POPKEY=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 CX POPUP= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

D

DMA/Audio-Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Drag & Drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

E

E-Mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

F

Fix Position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Frequency . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

G

GSE-Rapid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

H

Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 15 Hardwaretuning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 HARDWAREVERSION=. . . . . . . . . . . . . . . . . . . . . . . . . . 14 Hide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

I

icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Iconify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 IFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ILBM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 IR-Editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 IR-Runner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 IR-Runner Menus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 IR Appear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 IR Disable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 IR Disappear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 IR Move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 IR Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 IR Send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

J

JESOLO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 joystick port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Joystickport-Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

K

kHz. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Kickstart . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Kill Gadget(s) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Index 27

L

LANC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 LControl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 22 Learn IR data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Learnning IR-Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 LEFT= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Load backgound grafic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Load.... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 locale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

M

Make Screen Public . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Move Gadget(s). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

N

NewGadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Nico Francois . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20, 21

O

On public screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Other Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

P

Palette . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Parallel Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Parallelport-Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Prices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Publicscreen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 PUBSCREEN= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Pulse frequency . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Q

Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 12

R

Rapid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 22 Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Reichweite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Relais-Card . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 req.library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 reqtools.library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 RTPatch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

S

Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Save as... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Save Prefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Screenmode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Send command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Serial Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Show IR-Code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Size Gadge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Starting the IR-Runner . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

T

Test IR Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 The menus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 The Toolbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Toolbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Tooltypes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 TOP=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Tuning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Tuning (new Hardware). . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Tuning (old hardware) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

U

Update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Update hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

V

Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

W

What is IR-Master . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Wobbelcode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

Table of Contents

1 What is IR-Master . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 IR-Editor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1 The Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.1 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.2 Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.3 Arrange. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.4 Environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.1.5 Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.2 The Toolbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2.1 NewGadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2.2 Move Gadget(s) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2.3 Attache Grafic To Gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2.4 Attache Text to Gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2.5 Learn IR data. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2.6 Kill Gadget(s) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2.7 Size Gadge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2.8 Load Backgound Grafic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2.9 Change Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3.3 Test IR Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.4 Learning IR-Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.5 Drag & Drop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 IR-Runner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

4.1 Starting The IR-Runner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.2 IR-Runner Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3 ARexx Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.4 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

5 Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

5.1 Update old Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.2 Tuning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

5.2.1 Tuning the old Hardware. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 5.2.2 Tuning the new Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5.3 Joystickport-Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 5.4 Parallelport-Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.5 DMA/Audio-Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Appendix A Support & Bugreports . . . . . . . . . . . . . . . . . 18

A.1 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 A.2 Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

Appendix B Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Appendix C Other Software. . . . . . . . . . . . . . . . . . . . . . . . . . 20

C.1 req.library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 C.2 RTPatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 C.3 reqtools.library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 C.4 LControl & Rapid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 C.5 IR-Master History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26